### v2

# failed: manyfoot

failed: manyfoot
[…]
3]TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3, Canada
4]Laboratory of Instrumentation, Biomedical Engineering and Radiation Physics (LIBPhys-UNL), Department of Physics, NOVA School of Science and Technology, NOVA University Lisbon, 2829-516 Caparica, Portugal
5]Université de Strasbourg, CNRS, Institut de Physique et Chimie des Matériaux de Strasbourg,UMR 7504, F-67000 Strasbourg, France
6] Lawrence Livermore National Laboratory, Livermore, CA 94550, USA
7]STAR Cryoelectonics LLC, Santa Fe, NM 87508, USA
8]Department of Mettalurgical and Materials Engineering, Colorado School of Mines, Golden, CO 80401, USA
9]XIA LLC, Hayward, CA 94544, USA
10]LPC Caen, ENSICAEN, Université de Caen, CNRS/IN2P3, Caen, France
11]Department of Physics and Astronomy, McMaster University, Hamilton, Ontario L8S 4M1, Canada
12]Université Paris-Saclay, CEA, List, Laboratoire National Henri Becquerel (LNE-LNHB), F-91120, Palaiseau, France
13]Pacific Northwest National Laboratory, Richland, WA 99354, USA
Direct Experimental Constraints on the Spatial Extent of a Neutrino Wavepacket
Despite their high relative abundance in our Universe, neutrinos are the least understood fundamental particles of nature. They also provide a unique system to study quantum coherence in fundamental systems due to their extremely weak interaction probabilities. In fact, the quantum properties of neutrinos emitted in experimentally relevant sources are virtually unknown and theoretical predictions for the spatial width of neutrino wavepackets vary by many orders of magnitude. In weak nuclear decay, the size of a neutrino wavepacket, , is related to the spatial wavefunction of its parent at production. Experimentally, this value is only loosely constrained by neutrino oscillation data with a spread of 13 orders of magnitude. Here, we present the first direct limits of this quantity through a new experimental concept to extract the energy width, , of the recoil daughter nucleus emitted in the nuclear electron capture (EC) decay of Be. The final state in the EC decay process contains a recoiling Li nucleus and an electron neutrino () which are entangled at their creation. The Li energy spectrum is measured to high precision by directly embedding Be radioisotopes into a high resolution superconducting tunnel junction that is operated as a cryogenic charge sensitive detector. The lower limit on the spatial coherence of the recoil daughter was found to be pm, which implies the system remains in a spatially coherent state much larger than the nuclear scale. Further, this implies a lower limit on the size of a neutrino wavepacket, nm, which is more than five orders of magnitude more stringent than the limits from all combined reactor oscillation experiments. These results have wide-reaching implications in several areas including quantum coherence, the nature of spatial localization at sub-atomic scales, interpretation of neutrino physics data, and the potential reach of future large-scale experiments.
keywords:
quantum decoherence, neutrino, wavepacket, superconducting sensing, nuclear electron capture
Quantum mechanics is based on the concept that the microscopic scale of our Universe contains inherent uncertainties in position and momentum that are fundamentally connected via [1]. Using this theoretical description, precise knowledge of the momentum width of the particle, , implies a corresponding uncertainty on the extent of its spatial wavefunction, . When the object loses its ability to maintain coherence through interactions with the complex environmental bath its position is localized. This spatial localization of quantum objects depends heavily on the specific environment they are subjected to [2].
Recent efforts have primarily focused on constructing increasingly large coherent quantum systems [3, 4] and other highly delocalized objects to push measurement and other applications to, and beyond, the standard quantum limit [5]. Such experiments are at the frontiers of quantum science and engineering and aim, in part, to probe the interface between classical and quantum mechanics. Here, we present a new measurement concept on the other side of the coin; using highly localized systems with short coherence times embedded in a complex solid-state material at low temperature as a precision laboratory to probe quantum properties of particles by increasing their intrinsic energy or momentum broadening to a detectable level. Here, this method is shown to be particularly powerful for investigating properties of systems that are otherwise not directly accessible due to weak couplings in the standard model of particle physics (SM); in particular the neutrino ().
Neutrinos are light, neutral leptons, and the only particles in the SM that have an intrinsic chirality, in that they only interact via left-handed currents of the weak interaction through three eigenstates, , , and . Neutrino oscillation experiments over the last 30 years [6, 7] have also indicated that these flavour states include at least two non-zero mass eigenstates. This observation of non-zero neutrino masses currently provides the only confirmed violation of the SM as it was originally constructed [8, 9], for which the Nobel Prize in physics was awarded in 2015 [10]. The flavour oscillation effect for neutrinos arises from the fact that the mass eigenstates and weak-interaction eigenstates are not equal to each other, and due to their small interaction probabilities with the weak and gravitational forces, are able to maintain coherence over long distances. Thus, even for a neutrino created with a definite flavour – in weak nuclear decay, for example – a coherent superposition of states exists within the neutrino wavepacket that affects the probability for observing a particular flavour as a function of time or distance from the source.
The fact that neutrinos only interact weakly with matter makes them notoriously difficult to detect111For example, a with 1 MeV kinetic energy has an interaction cross section of roughly cm [11]. This fact also makes neutrinos attractive systems to study fundamental properties of quantum mechanics and exotic new physics [12]. Historically, observing these so-called “ghost particles” from various sources requires large volume detectors with limited detection precision that run for several years [13]. These detectors rely on neutral current (NC) and charged current (CC) interactions with an atomic nucleus or electron within these large volumes. The final state particles are used to reconstruct the incoming neutrino energy and interaction point, along with neutrino flavour for CC interactions. A deficit or excess flux of any flavour in the detector, compared to the flavour ratios at the source, is used to determine the oscillation probabilities as a function of energy and distance. Global data from neutrino oscillation experiments have been used to determine the mass-splittings and mixing parameters for the three SM neutrinos [14]. Data from the Daya Bay [15, 16], RENO [17, 18], and KamLAND [19] experiments currently also provide loose constraints on the spatial extent of wavepackets from decay (reactor sources) of m m [20, 21, 22]. If the and wavepacket spatial widths are indeed near the low end of these limits, it has a significant impact on the neutrino physics landscape, including alleviating tension in models for sterile neutrinos with eV-scale masses through model-dependent fits to available data [23, 24].
To extract physics from the oscillation data, however, interpretation within the lepton mixing infrastructure of the SM is required [14]. Given the current pervading anomalies in the neutrino physics landscape [25], there is a need for new experimental techniques that can provide model-independent access to the neutrino. Clever new experimental concepts have been presented to address this need using efficient, small scale setups [26, 27, 28] that are less model-dependent and not limited by low neutrino interaction cross sections. These experiments aim to precisely measure the low-energy recoiling atoms in nuclear electron capture (EC) decay [29] to access information of the neutrino directly through its entanglement in the final state of radioactive decay. Of these new concepts, the beryllium electron capture in superconducting tunnel junctions (BeEST) experiment [26] is currently the only one to practically employ the concept of direct EC daughter recoil detection for neutrino physics [30] and astrophysics [31]. In this paper, we report the first direct limit on the energy width of the recoiling nucleus in EC decay, , to elucidate the spatial width of a neutrino wavepacket through their mutual entanglement. Using the high-resolution BeEST data, we report an improvement on the lower limit of the size of a neutrino wavepacket by over 5 orders of magnitude and comment on the quantum localization scale of weak nuclear decay.
Nuclear EC decay is a weak-interaction-induced radioactive decay mode that is analogous to nuclear decay, but involves the capture of an orbital atomic electron in the initial state and generates a final state that, to first order, only includes two products: the recoiling transmuted daughter atom and [29]. This simple two-body system is entangled at its creation, and thus a precision measurement of the daughter recoil provides direct access to information on the neutrino.
Be decays by EC primarily to the nuclear ground-state of Li with a half-life of days [32]. A small branch of 10.44(4)% results in the population of a short-lived excited nuclear state in Li ( fs) [32] that de-excites via emission of a 477.603(2) keV -ray [33]. In the EC process, the electron can be captured either from the shell (-capture) or the shell (-capture) of Be. For -capture, the binding energy of the hole is subsequently liberated by emission of an Auger electron [34] whose energy adds to the decay signal and separates it from the -capture signal. Since the nuclear decay and subsequent atomic relaxation occur on short time scales, a direct measurement produces a spectrum with four peaks: two for -capture and two for -capture into the ground state (GS) and the excited state (ES) of Li.
Figure 1: A schematic image of the STJ sensor showing the five layers described in the text. The Be radioactive source is sparsely implanted into the Ta lattice of the absorber layer. The upper right zoom-in shows the final state after EC decay including the Li recoil and . The Li recoil inherits the spatial uncertainty from its parent nucleus which was embedded in the Ta lattice, and sets the scale for the width of the emitted wavepacket.
To measure the Be EC decay recoil broadening to high precision, we directly embedded the Be into the top tantalum layer of a thin superconducting tunnel junction (STJ) pixel that was used as a high energy-resolution sensor [35]. The single STJ pixel used in this experiment was part of a 36-pixel array where each sensor had a surface area of m and consisted of five layers (from top to bottom): Ta (165 nm)—Al (50 nm)—AlO (1 nm)—Al (50 nm)—Ta (265 nm), shown in Fig. 1, that were fabricated by photolithography at STAR Cryoelectronics LLC [36]. STJs exploit the small energy gap in superconducting Ta ( meV) to provide higher energy resolution than conventional Si or Ge detectors, thus allowing for a precision measurement of the Li recoil energies.
The radioactive Be ions were implanted directly into the tantalum layer at the top of the STJ pixel through apertures in a Si collimator that was situated m above the chip. The Be ions were produced and delivered to the implantation chamber end-station at the TRIUMF-ISAC facility in Vancouver, Canada [37] at an energy of 30 keV. The isotopically pure Be beam was produced using the isotope separation on-line (ISOL) technique [38] via spallation reactions from a 10 A, 480-MeV proton beam incident on a stack of thin uranium carbide targets. Once released from the target through diffusion, the created beryllium atoms were selectively ionized using the ion guide laser ion source (IG-LIS) [39] operated in suppression mode [40]. The resulting beam rate of Be ions at the implantation end station was and was implanted for a period of 25 hours generating an initial per-pixel activity in the STJs of roughly 50 Bq. Following implantation, the chip was cleaned and processed at TRIUMF to remove loose and surface-deposited activity and subsequently shipped to Lawrence Livermore National Laboratory (LLNL).
The Li recoil spectrum from the decay of Be was measured at LLNL with the STJ detector at a temperature of 0.1 K in a wet two-stage adiabatic demagnetization refrigerator (ADR) with liquid nitrogen and liquid helium precooling. Signal traces were read out continuously at a rate of 1.25 MSa/s at a 16 bit resolution using a PXIe-6356 ADC. Our data acquisition system achieved a preamplifier-limited resolution of eV. For real-time in-situ monitoring of the response and energy calibration, the STJs were simultaneously exposed to 3.49865(15) eV photons from a pulsed, 355 nm frequency tripled Nd:YVO laser triggered at a rate of 100 Hz that was fed into the cold stage of the ADR through a fiber to illuminate the pixels. The laser intensity was adjusted such that multi-photon absorption provided a comb of peaks over the energy range from 20-120 eV.
Data were acquired for 20 hours on November 6, 2022. Voltage traces from the STJ array were continuously recorded along with the laser and processed using a trapezoidal filter. The pulsed laser was used as a near event-by-event energy calibration and drift correction over the period of data collection. The laser events were selectively identified in the data stream via their coincidence with a logic pulse that was used to trigger the laser, and the comb of peaks is shown in Fig. 2. The laser signal was fit to a superposition of individual Gaussian functions, each corresponding to an integer multiple of the single photon energy to generate a quadratic calibration function which was applied to the corresponding Be spectrum. Using this approach, we achieved a precision on the energy residuals of the calibration 0.01 eV in the region of interest below 200 eV [41, 42].
Figure 2: The measured Li recoil spectrum with the four peaks described in the text for data from a single STJ pixel shown in the inset. The measured of the K-GS peak is shown, and is conservatively extracted as the upper limit on the energy width of the recoil. The calibration laser spectrum (violet) is also shown for comparison.
The final Be decay spectrum shown in Fig. 2 includes events from a single pixel that were not in coincidence with the laser logic pulse or simultaneous signals in other pixels. The latter of these two criteria was imposed to remove the large, low-energy background from 478 keV -ray Compton scattering interactions in the 0.5 mm Si substrate below the STJ array which generated signals in many pixels simultaneously. The resulting spectrum contains four main peaks resulting from the two atomic and two nuclear processes described above. The two recoil peaks from the excited state decay of Li were Doppler-broadened from -decay in flight [31, 30].
Since the most probable Be decay mode is capture to the nuclear ground state of Li [31], the K-GS peak provides the highest statistical significance for a measurement of the energy width. Since the Be is sequestered within the superconducting Ta lattice, the wavefunctions of the Be orbital electrons are hybridized with the Ta lattice (and impurities) which contribute of the broadening measured in the K-GS peak [43], which is likely the dominant source for the measured peak broadening. However, due to their complexity and the fact that these effects are not yet quantified in detail for the BeEST energy spectrum, we conservatively attribute all of the measured broadening in the K-GS peak to quantum uncertainty. To extract the energy width, the K-GS peak in the single-pixel energy spectrum shown in the inset of Fig. 2 was fit in the region from 100-111 eV with a single Voigt function [44, 45]. We then set an upper limit on the uncertainty in the nuclear recoil energy measurement at 2.9 eV with a 95% confidence level (CL) using the profile likelihood method [46].
The limits on were extracted using two approaches [47, 48]. The first assumes that the neutrino energy width is equivalent to that of the entangled recoil nucleus, , and these widths are determined by the timescale of interactions within the production environment. This leads to a velocity dependent relation between the spatial width and energy width [47]:
In lieu of any previous measurements of , Refs. [49, 50] estimate the size of the neutrino wavepacket with assumptions for the average time between atomic interactions of the parent and decay products within a material which results in . Here, our measurement of eV requires no assumption of scale under this approach, and provides direct limits on the size of the neutrino wavepacket of nm at 95% CL, consistent with the above estimates.
The second approach [48] considers momentum space wavepackets produced in a two-body decay leading to the and nuclear position uncertainties being related by:
where, for the BeEST experiment, and are the atomic masses of Be and Li [51] respectively, and is the initial momentum of the Be. In Ref. [48], the assumption is made that the nuclear wavepacket is localized on the scale of nucleon or nuclear diameters (1-6 fm), and not the atomic interaction scale, resulting in estimates for from decays in nuclear reactors of pm. To compare our results with this approach, we convert our direct measurement of nuclear recoil energy uncertainty to a momentum uncertainty using the classical relation . This can be done by either propagating the uncertainty or by converting the energy distribution to a momentum distribution and fitting this to a Voigt function. Both result in the same value at the 95% CL. We then relate this momentum uncertainty to the nuclear position uncertainty through the Heisenberg uncertainty principle,
Under this approach, we obtain pm and nm at 95% CL, which is inconsistent with the assumption that the weak decay localization scale of the daughter is on the nuclear scale [48].
Figure 3: Experimental lower limits at 95% CL (arrows) and theoretical predictions based on the quantum mechanical localization scale of the decay (solid bars) for the spatial extent of a neutrino wavepacket, , generated by a radioactive decay source. The top line shows the BeEST limit (this work) and the following two lines show experimental limits from the reactor data in the Daya Bay experiment [20] and a combined analysis of the Daya Bay, RENO, and KamLAND experimental data [22]. The coloured regions denote predicted ranges for neutrino wavepacket widths for assumptions that the localization scale of weak nuclear decay is on the nuclear/nucleon scale [48] or the atomic/interatomic scale [49, 50]. The Daya Bay collaboration also places an upper limit of m based on the size of their detector and source that is not shown in this plot [20].
Since both methods described above are discussed extensively in the literature [49, 52, 53, 48], we do not claim validity of one over the other here, but rather use our results from the more conservative of the two to report a lower limit of nm at 95% CL. When compared to all previous data from reactor source oscillation experiments, we improve the experimental lower limit on the size of a nuclear decay induced neutrino wavepacket, , by more than five orders of magnitude (Fig. 3). Assuming any potential variation in between and EC decay is much smaller than the variation caused by the localization scale of the source, the presented limit implies that neutrino wavepacket separation is not causing the dampening-like effect preferred by the 3+1 sterile neutrino () model fits [23, 24], and thus these anomalies still persist. Our results here also suggest that concerns of oscillation damping in large-scale experiments such as JUNO [54] due to wavepacket separation will not be visible, and thus have no effect on the planned program.
Beyond the implications for neutrino physics, perhaps the more important result from our work is the potential to use weak nuclear decay as a fundamental test of quantum mechanics [48]. Since the neutrino is produced within the nuclear volume via weak coupling at the quark level, there is a question as to whether processes such as short-range correlations [55] are able to decohere and localize the system via interactions within the atomic nucleus (10 fm) [48]. Our experimental determination of the position width of the daughter recoil, pm, excludes this possibility by several orders of magnitude as illustrated in Fig. 3. This general conclusion is consistent with nuclear decay, an electromagnetic process, in highly localized environments through the Mössbauer effect, which suggests localization on the scale of 1-10 pm [56].
In summary, we present a new experimental concept of using highly localized rare isotopes inside superconducting sensors to measure the spatial extent of a neutrino wavepacket. Our measurement provides the first direct limits on this quantity through a precision measurement of the Li daughter recoil energy width produced in the EC decay of Be. The lower limit nm at 95% CL is 5 orders of magnitude more stringent than those from all combined reactor oscillation data. This result implies that the neutrino wavepacket from radioactive sources is too large to observe wavepacket separation in future large scale experiments and does not relieve the tension in 3+1 model fits. The inferred spatial width of the daughter following EC, pm, excludes the possibility that quantum localization occurs on the nuclear scale. These results have implications for the details of neutrino oscillation and BSM theories. They may also suggest that the neutrino emitted from weak nuclear decay in a solid-state lattice may be large enough to be detected with macroscopic coherent quantum objects or exploited in other ways for applied neutrino science.
Acknowledgements
We thank David Moore, Daniel Carney, Benjamin Jones, Carlos Argüelles-Delgado, Joseph Formaggio, and Fred Sarazin for useful discussions. The BeEST experiment is funded by the Gordon and Betty Moore Foundation (10.37807/GBMF11571), the U.S. Department of Energy Office of Science, Office of Nuclear Physics under Award Numbers DE-SC0021245 and SCW1758, the LLNL Laboratory Directed Research and Development program through Grants No. 19-FS-027 and No. 20-LW-006, the European Metrology Programme for Innovation and Research (EMPIR) Projects No. 17FUN02 MetroMMC and No. 20FUN09 PrimA-LTD, and the FCT—Fundação para a Ciência e Tecnologia (Portugal) through national funds in the framework of the projects UID/04559/2020 (LIBPhys). TRIUMF receives federal funding via a contribution agreement with the National Research Council of Canada. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344. Francisco Ponce is supported by Pacific Northwest National Laboratory, which is managed for the US Department of Energy by Battelle under contract DE-AC05-76RL01830. KGL gratefully acknowledges support from the Facility for Rare Isotope Beams (FRIB) while on sabbatical. FRIB is a US Department of Energy, Office of Science User Facility under Award Number DE-SC0000661.